Assume that Hank and Mika can switch between producing baseballs and producing bats at a constant rate.

Labor Hours Needed
to Make 1

Quantity Produced
in 36 Hours

baseball

bat

baseballs

bats

Hank

2

6

18

6

Mika

2

4

18

9



Refer to Table 3-8. Mika has an absolute advantage in the production of

a.

bats and a comparative advantage in the production of baseballs.

b.

bats and a comparative advantage in the production of bats.

c.

neither good and a comparative advantage in the production of baseballs.

d.

baseballs and a comparative advantage in the production of bats.
Answer #1

For Mika

Opportunity cost to produce bat = 18 / 9 = 2 baseballs

Opportunity cost to produce baseballs = 9 / 18 = 0.5 bats

For Hank

Opportunity cost to produce bat = 18 / 6 = 3 baseballs

Opportunity cost to produce baseballs = 6 / 18 = 0.333 bats

So Mika has an absolute advantage in production for bats and comparative advantage in the production of bats

So correct answer is option b. bats and a comparative advantage in the production of bats.